OpenStudy (anonymous):

find the maximum and minimum of: y= 7+3x-2x^(2)

OpenStudy (anonymous):

got a clue how to the max and min?

OpenStudy (anonymous):

yeah.... first is to know its first derivative which is: y'=-4x + 3 but i don't know how can i get its' critical values...

OpenStudy (anonymous):

critical points is when you set the derivative equal to 0 y'=0=-4x+3 solve for x and find y(x) just by looking at the function, we know its a parabolic curve, meaning there's either a maximum or a minimum, unless you have boundries so you just need to determine whether (x,y(x)) is a maximum or a minimum

OpenStudy (anonymous):

i meant parabola

OpenStudy (anonymous):

so.... x=3/4 then, y= 65/8 but how to find the minima & maxima?

OpenStudy (anonymous):

is it necessarily to have two critical values?

OpenStudy (anonymous):

if you had x^3+..., then you would have 2 critical values, but for parabolas, theres only 1 max/min

OpenStudy (anonymous):